OpenStudy (anonymous):

I WILL BECOME A FAN N GIVE A MEDAL!!!<<<<<<< Danny uses an app that shows him how many kilometers he has run to prepare for a marathon. The app said he ran 8.045 kilometers. He wants to post online how many miles he ran. Danny ran ___ miles. (1 mile = 1.609 kilometers) Numerical Answers Expected! Answer for Blank 1:

OpenStudy (anonymous):

can i please get help with this WARNING:AWARD MEDAL,AND I WILL BECOME A FAN

OpenStudy (gorv):

1.609km=1mile so 1km=1/1.609 mile

OpenStudy (anonymous):

so is the answer 1.609 @gorv

OpenStudy (gorv):

now 8.045km=8.045/1.609

OpenStudy (gorv):

\[1km=\frac{ 1 }{ 1.609 }\]

OpenStudy (anonymous):

can someone please tell me if 1.609 is the answer?